OpenStudy (anonymous):

Hey, if 2|x+5|=10 and 4|3-y|=16, then x+y could equal each of the following except (a) -11 (b) -3 (c) -1 (d) 7

OpenStudy (blockcolder):

Use the Reverse Triangle Inequality: \[\left|a-b\right|\geq\left|(|a|-|b|)\right|\] And let a=x+5, b=3-y.